Operations Report – January 9th

Crew 202 Operations Report 09-JAN-2019

SOL: 11

Name of person filing report: Kasey Hilton

Non-nominal systems: Hab ceiling leak; Suit #7 battery

Notes on non-nominal systems: The Hab ceiling is leaking over the couch area this morning but stopped in the afternoon, most likely due to the extra moisture from the melting snow; The battery in suit #7 was very low after the EVA and the fans were barely blowing; The suit is going to charge overnight and I will check the battery voltage again in the morning

Generator (hours run): 17hr 3min; Turned on last night (08Jan2019) at 16:33; Turned off this morning (09Jan2019) at 9:36; Turned on tonight (09Jan2019) at 15:15

Solar SOC – Turned on (08Jan2019) 80%; Turned off (09Jan2019) 100%; Turned on (09Jan2019) 82%

Diesel Reading – 70%

Propane Reading – 37%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 50%; 270 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 60%; 178 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 18:28 to refill the tank

Water Meter: 01399232

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes

Hours: 66.8

Beginning charge: 100%

Ending charge: 26%

Currently charging: Yes

Opportunity rover used: No, brakes are still not working correctly

Hours: 45.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 68.7

Beginning charge: 100%

Ending charge: 45%

Currently charging: Yes

Notes on rovers: Brake fluid in Opportunity was replaced but the brakes are still not functioning properly

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: Nothing to report

HabCar used and why, where? Not used

CrewCar used and why, where? Off-site

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Battery in suit #7 was very low by the end of the EVA as mentioned above in “Non-nominal systems”; Suit will be charged overnight and battery voltage will be checked again in the morning

Summary of Hab operations: Hab ceiling leaking in one spot in the corner as mentioned above in “Non-nominal systems”

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Hot plate was cleaned and will be tested tomorrow to see if the odor continues

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing to report